# Cosmological Stasis from a Single Annihilating Particle Species: Extending Stasis Into the Thermal Domain

### v1

@@ Abstract
It has recently been shown that extended cosmological epochs can exist during which the abundances associated with different energy components remain constant despite cosmological expansion. Indeed, this “stasis” behavior has been found to arise generically in many BSM theories containing large towers of states, and even serves as a cosmological attractor. However, all previous studies of stasis took place within non-thermal environments, or more specifically within environments in which thermal effects played no essential role in realizing or sustaining the stasis. In this paper, we demonstrate that stasis can emerge and serve as an attractor even within thermal environments, with thermal effects playing a critical role in the stasis dynamics. Moreover, within such environments, we find that no towers of states are needed — a single state experiencing two-body annihilations will suffice. This work thus extends the stasis phenomenon into the thermal domain and demonstrates that thermal effects can also generally give rise to an extended stasis epoch, even when only a single matter species is involved.
@@ I Introduction
It has recently been observed that many theories of physics beyond the Standard Model (BSM) give rise to early-universe cosmologies which potentially contain extended epochs during which the abundances of different cosmological energy components (such as matter, radiation, and vacuum energy) remain constant despite cosmological expansion. This phenomenon has been dubbed “stasis” Dienes et al. (2022a), and seems to be rather ubiquitous in systems containing large (or even infinite) towers of increasingly massive states, such as are known to arise in many BSM models Barrow et al. (1991); Dienes et al. (2022a, b, 2023, 2024); Halverson and Pandya (2024). Indeed, it has been found that matter and radiation can be in stasis with each other Dienes et al. (2022a); Barrow et al. (1991); Dienes et al. (2022b); Halverson and Pandya (2024), and that each of these can also be in stasis with vacuum energy Dienes et al. (2023, 2024). It has even been found that matter, radiation, and vacuum energy can all experience a simultaneous triple stasis Dienes et al. (2023). Moreover, stasis is a dynamical attractor within such cosmologies. Thus, even if these systems do not begin in stasis, they will naturally flow towards such stasis configurations. This renders the stasis state essentially unavoidable in many BSM cosmologies.
@@
All of the BSM cosmologies which have been examined thus far share certain characteristics. First, they are all non-thermal. By this, we mean that they do not involve any cosmological energy components whose temperatures play a critical role in the stasis phenomenon. Even in cases for which the stasis is realized through the Hawking radiation emitted by primordial black holes Barrow et al. (1991); Dienes et al. (2022b), it is sufficient to regard the black-hole evaporation process as one which simply converts matter (in the form of black holes) to radiation. In such analyses — indeed, in all of the different stasis scenarios examined thus far — the temperatures that might happen to characterize particular cosmological populations of particles play no essential role in the stasis dynamics. Secondly, and perhaps even more importantly, in all of the BSM systems examined thus far the existence of a bona fide tower of different species $\phi_{\ell}$ with different masses $m_{\ell}$ was critical. Indeed, we have even repeatedly found that the duration of the resulting stasis epoch is directly related to the number of species in the tower. Finally — and a bit more technically — the energy-transfer mechanisms that drove the stasis phenomenon in each of the previous cases (such as the decay of matter to radiation in Refs. Dienes et al. (2022a); Barrow et al. (1991); Dienes et al. (2023), or the transition from overdamped fields to underdamped fields in Refs. Dienes et al. (2023, 2024)) had the property that the rate of energy transfer depended linearly on the energy density associated with the original field. More specifically, the “pumps” $P^{(\rho)}$ that were required in each of the previous cases were each proportional to $\rho$. This linearity property played a significant role in establishing the algebraic structure of the resulting stasis. While suitable for decay processes in which a matter particle $\phi$ decays into radiation, thereby leading to a rate of energy transfer from matter to radiation scaling linearly with the $\phi$-particle density $n_{\phi}$, this linearity property would seem to preclude processes such as particle annihilation, for which we would expect a rate that scales as $n_{\phi}^{2}$.
@@
As might be imagined, the thermal stasis that we shall describe in this paper is quite different from the non-thermal stases that already exist in the literature. Without a tower of states, and with the stasis driven by particle annihilation rather than by particles decaying or undergoing phase transitions, it might seem at first glance that thermal stasis must emerge in a completely different way. Indeed, even the pumps that we shall employ in this paper may be more familiar from studies of thermal freezeout in weakly-interacting-massive-particle (WIMP) models of dark matter than from prior studies of stasis. However, despite these important differences, we shall nevertheless find that a stasis likewise emerges within this new context and shares many critical properties with its non-thermal cousins.
@@ II Assumptions and Energy Components
Let us start by considering the cosmology associated with a universe which comprises a non-relativistic matter species $\phi$ of mass $m$ which annihilates into an effectively massless particle $\chi$ through processes of the form $\phi\phi\to\chi\chi$. For simplicity, we shall assume that $\phi$ and $\chi$ are both their own antiparticles (although our qualitative results would be unchanged if we were to drop this last assumption). Thus, such a universe contains not only a matter component $\phi$ with an energy density $\rho_{M}$, abundance $\Omega_{M}$, and equation-of-state parameter $w_{M}=0$, but also a radiation component $\chi$ with a corresponding energy density $\rho_{\gamma}$, abundance $\Omega_{\gamma}$, and equation-of-state parameter $w_{\gamma}=1/3$. For simplicity, we shall assume that these matter and radiation components are the only cosmological energy components present within our universe, and that the universe is flat and expanding in a manner consistent with the FRW metric.
Remarkably, we shall demonstrate that such a universe can give rise to an extended stasis epoch. To do this, we shall assume that our population of matter particles comprises only a single species $\phi$ with a unique mass $m$. We shall also assume that this species is stable in the sense that it does not decay within the time frame relevant for our analysis. Finally, we shall assume that the $\phi$ particles are not absolutely cold, but rather are in thermal equilibrium with each other, possibly as the result of rapid interactions among them, and constitute an ideal gas with a non-zero temperature $T$.
@@
Given these assumptions, we see that we must actually regard our universe as containing three different “fluids”. The first two fluids are those mentioned above, namely those associated with matter and radiation, but our matter fluid really only accounts for the rest-mass energy of our matter fields. We must therefore now introduce a third fluid, namely that associated with the kinetic energy of our matter fields, with corresponding energy density $\rho_{\rm KE}$, abundance $\Omega_{\rm KE}$, and equation-of-state parameter $w_{\rm KE}$. It is important to note that we cannot consider $\rho_{M}$ and $\rho_{\rm KE}$ as different contributors to a single common fluid associated with our gas of $\phi$ particles with a unique time-independent equation of state of the form $P=w\rho$ with a constant $w$, since (as we shall demonstrate below) the energy densities associated with rest mass and kinetic energy have different equations of state of this form, with $w_{M}\not=w_{\rm KE}$.
@@ III Dynamical Equations
In general, following Ref. Dienes et al. (2023), we shall refer to any process that induces a transfer of energy between two energy components as a “pump.” Such pumps are associated with local particle-physics processes which conserve energy density, but redistribute it among our different energy components.
@@ IV Pump Terms for Thermal Annihilation
It is not difficult to obtain explicit expressions for our pump terms. Within the cosmology we are studying, such pumps represent the effects of the two-body annihilations of $\phi$ particles into radiation. Such process are familiar from traditional studies of the thermal freezeout phenomenon, where one has $dn_{M}/dt\sim\langle\sigma v\rangle n_{M}^{2}+...$ where $n_{M}$ is the matter-particle number density. In this expression, $\langle\sigma v\rangle$ denotes the thermally averaged “swept-volume” rate, i.e., the rate at which volume is swept by the cross-section $\sigma$ moving with transverse velocity $v$. Recognizing that $n_{M}=\rho_{M}/m$, we see that our corresponding pump is nothing but
@@
where $|\vec{p}_{\rm CM}|\equiv|\vec{p}_{a}-\vec{p}_{b}|$ is the magnitude of the three-momentum of either of the two incoming particles $\phi_{a}$ and $\phi_{b}$ in the CM frame, where $q$ is an arbitrary exponent, and where $C$ is a momentum-independent and $q$-independent coefficient. The ansatz in Eq. (16) is not entirely unfamiliar; for example, a similar ansatz is often invoked when discussing thermal freezeout (see, e.g., Ref. Kolb and Turner (1990); Jungman et al. (1996); Bertone et al. (2005)), with the exponent $q$ selecting between $s$-wave annihilation ($q=0$), $p$-wave annihilation ($q=2$), $d$-wave annihilation ($q=4$), and so forth. However, although the value of $q$ naïvely corresponds to twice the order of the leading term in the partial-wave expansion, $q$ can actually be more negative than stated above if our annihilation proceeds through an $s$-channel process with a propagator that is in resonance at low momenta. Furthermore, Sommerfeld enhancement Sommerfeld (1931) can also give rise to negative values of $q$. For this reason, we will also allow ourselves to consider values of $q$ which are negative. Indeed, we shall give an example of a model that satisfies Eq. (16) with such values of $q$ in Sect. VIII.
@@ V Dynamical Equations Redux and the Role of Coldness
Given these pumps, we can now construct the differential equations that govern the time-evolution of our system within this cosmology. In principle, there are two quantities whose variations are of interest to us: $\Omega_{M}$ and $\Omega_{\rm KE}$. Indeed, as is evident from Eq. (2), the temperature $T$ is simply related to the quotient $\Omega_{\rm KE}/\Omega_{M}$. However, $\rho_{M}$ remains as an additional independent quantity within this system because the expressions in Eq. (25) for our energy-density pumps depend quadratically on $\rho_{M}$. Indeed, although one of these factors of $\rho_{M}$ is converted to $\Omega_{M}$ when constructing our abundance pumps $P_{ij}$ in Eq. (10), the other factor of $\rho_{M}$ remains unconverted. Thus, in this system, our differential equations cannot be reduced to abundances alone, and the absolute matter energy density $\rho_{M}$ — or equivalently the Hubble parameter — remains as an additional degree of freedom. This is a new feature that did not appear in any of the previous stasis analyses in Refs. Dienes et al. (2022a, b, 2023). However, this feature arises in the present case as a new consequence of our choice of pump.
@@ VI Fixed-Point Solutions and Thermal Stasis
Interestingly, we see that the fixed-point abundance ${\overline{\Omega}}_{M}$ does not depend on the pump prefactor $\widehat{C}$. This is ultimately the case because the dynamical equations in Eq. (28) are invariant under the simultaneous transformations $\Xi\to\alpha\Xi$ and $\widehat{C}\to\widehat{C}/\sqrt{\alpha}$, where $\alpha$ is an arbitrary scaling parameter. This observation is analogous to the observation that the quantity ${\overline{\Omega}}_{M}$ in Ref. Dienes et al. (2022a) is independent of $\Gamma_{0}$, as well as similar observations in Refs. Dienes et al. (2022b, 2023).
@@
We thus conclude that our system necessarily evolves into an extended stasis epoch, regardless of its initial conditions pertaining to either initial matter (or radiation) abundance or initial coldness. This is therefore a direct demonstration that the stasis phenomenon can be extended into the thermal domain, with a prolonged period of stasis experienced not only by the energy abundances but also by thermodynamic quantities — such as coldness — which involve the temperature of the matter prior to annihilation.
@@
Given this state of affairs, it is natural to wonder how this stasis is able to persist across so many $e$-folds of cosmological expansion — a concern heightened by the fact that only a single matter species is involved. Ultimately, this feature is tied to the unique form of our pump. As we have seen, our $\phi$-particle annihilation pumps for both energy density and kinetic energy density scale as
@@ VII Thermal Stasis as a Global Attractor
It is straightforward to demonstrate that the non-trivial fixed-point solutions for $({\overline{\Omega}}_{M},{\overline{\Xi}})$ in Eq. (32) are also global attractors. We have already seen evidence of this attractor behavior in Fig. 2. As a first step, we note that for a given value of $q$, we may assess whether that the corresponding fixed-point solution is a local attractor by evaluating the $2\times 2$ Jacobian matrix for our dynamical equations for $d\Omega_{M}/d{\cal N}$ and $d\Xi/d{\cal N}$ and then examining the signs of its two eigenvalues $\lambda_{\pm}$ at that fixed point. In Fig. 3, we present a parametric plot of these eigenvalues as $q$ is varied within the range $q_{\rm min}<q<q_{\rm max}$ for which our non-trivial fixed-point solution exists. These eigenvalues trace out the red solid curve. Within this $q$-range, we see that both eigenvalues lie within the yellow-shaded region wherein both $\lambda_{+}$ and $\lambda_{-}$ are negative. Thus, we may conclude that across this entire range of $q$, the non-trivial fixed-point solution within this range is indeed a local attractor.

### v2

@@ Abstract
It has recently been shown that extended cosmological epochs can exist during which the abundances associated with different energy components remain constant despite cosmological expansion. Indeed, this “stasis” behavior has been found to arise generically in many beyond-the-Standard-Model theories containing large towers of states, and even serves as a cosmological attractor. However, all previous studies of stasis took place within non-thermal environments, or more specifically within environments in which thermal effects played no essential role in realizing or sustaining the stasis. In this paper, we demonstrate that stasis can emerge and serve as an attractor even within thermal environments, with thermal effects playing a critical role in the stasis dynamics. Moreover, within such environments, we find that no towers of states are needed — a single state experiencing two-body annihilations will suffice. This work thus extends the stasis phenomenon into the thermal domain and demonstrates that thermal effects can also generally give rise to an extended stasis epoch, even when only a single matter species is involved.
@@ I Introduction
It has recently been observed that many theories of physics beyond the Standard Model (BSM) give rise to early-universe cosmologies which potentially contain extended epochs during which the abundances of different cosmological energy components (such as matter, radiation, and vacuum energy) remain constant despite cosmological expansion. This phenomenon has been dubbed “stasis” Dienes et al. (2022a), and seems to be rather ubiquitous in systems containing large (or even infinite) towers of increasingly massive states, such as are known to arise in many BSM models Barrow et al. (1991); Dienes et al. (2022a, b, 2024a, 2024b); Halverson and Pandya (2024). Indeed, it has been found that matter and radiation can be in stasis with each other Dienes et al. (2022a); Barrow et al. (1991); Dienes et al. (2022b); Halverson and Pandya (2024), and that each of these can also be in stasis with vacuum energy Dienes et al. (2024a, b). It has even been found that matter, radiation, and vacuum energy can all experience a simultaneous triple stasis Dienes et al. (2024a). Moreover, stasis is a dynamical attractor within such cosmologies. Thus, even if these systems do not begin in stasis, they will naturally flow towards such stasis configurations. This renders the stasis state essentially unavoidable in many BSM cosmologies.
@@
All of the BSM cosmologies which have been examined thus far share certain characteristics. First, they are all non-thermal. By this, we mean that they do not involve any cosmological energy components whose temperatures play a critical role in the stasis phenomenon. Even in cases for which the stasis is realized through the Hawking radiation emitted by primordial black holes Barrow et al. (1991); Dienes et al. (2022b), it is sufficient to regard the black-hole evaporation process as one which simply converts matter (in the form of black holes) to radiation. In such analyses — indeed, in all of the different stasis scenarios examined thus far — the temperatures that might happen to characterize particular cosmological populations of particles play no essential role in the stasis dynamics. Secondly, and perhaps even more importantly, in all of the BSM systems examined thus far the existence of a bona fide tower of different species $\phi_{\ell}$ with different masses $m_{\ell}$ was critical. Indeed, we have even repeatedly found that the duration of the resulting stasis epoch is directly related to the number of species in the tower. Finally — and a bit more technically — the energy-transfer mechanisms that drove the stasis phenomenon in each of the previous cases (such as the decay of matter to radiation in Refs. Dienes et al. (2022a); Barrow et al. (1991); Dienes et al. (2024a), or the transition from overdamped fields to underdamped fields in Refs. Dienes et al. (2024a, b)) had the property that the rate of energy transfer depended linearly on the energy density associated with the original field. More specifically, the “pumps” $P^{(\rho)}$ that were required in each of the previous cases were each proportional to the corresponding energy density $\rho$. This linearity property played a significant role in establishing the algebraic structure of the resulting stasis. While suitable for decay processes in which a matter particle $\phi$ decays into radiation, thereby leading to a rate of energy transfer from matter to radiation scaling linearly with the $\phi$-particle density $n_{\phi}$, this linearity property would seem to preclude processes such as particle annihilation, for which we would expect a rate that scales as $n_{\phi}^{2}$.
@@
As might be imagined, the thermal stasis that we shall describe in this paper is quite different from the non-thermal stases that already exist in the literature. Without a tower of states, and with the stasis driven by particle annihilation rather than by particles decaying or undergoing phase transitions, it might seem at first glance that thermal stasis must emerge in a completely different way. Indeed, even the pumps that we shall employ in this paper may be more familiar from studies of thermal freeze-out in weakly-interacting-massive-particle (WIMP) models of dark matter than from prior studies of stasis. However, despite these important differences, we shall nevertheless find that a stasis likewise emerges within this new context and shares many critical properties with its non-thermal cousins.
@@ II Assumptions and Energy Components
Let us start by considering the cosmology associated with a universe which comprises a non-relativistic matter species $\phi$ of mass $m$ which annihilates into an effectively massless particle $\chi$ through processes of the form $\phi\phi\to\chi\chi$. For simplicity, we shall assume that $\phi$ and $\chi$ are both their own antiparticles (although our qualitative results would be unchanged if we were to drop this last assumption). Thus, such a universe contains not only a matter component $\phi$ with an energy density $\rho_{M}$, abundance $\Omega_{M}$, and equation-of-state parameter $w_{M}=0$, but also a radiation component $\chi$ with a corresponding energy density $\rho_{\gamma}$, abundance $\Omega_{\gamma}$, and equation-of-state parameter $w_{\gamma}=1/3$. For simplicity, we shall assume that $\phi$ and $\chi$ are the only cosmological species present within our universe, and that the universe is flat and expanding in a manner consistent with the Friedmann-Robertson-Walker (FRW) metric.
Remarkably, we shall demonstrate that such a universe can give rise to an extended stasis epoch. To do this, we shall assume that our population of matter particles comprises only a single species $\phi$ with a unique mass $m$, as noted above. We shall also assume that this species is stable in the sense that it does not decay within the time frame relevant for our analysis. Finally, we shall assume that the $\phi$ particles are not absolutely cold, but rather are in thermal equilibrium with each other, possibly as the result of rapid interactions among them, and constitute an ideal gas with a non-zero temperature $T$.
@@
Given these assumptions, we see that we must actually regard our universe as containing three different “fluids.” The first two fluids are those mentioned above, namely those associated with matter and radiation, but our matter fluid really only accounts for the rest-mass energy of our matter fields. We must therefore now introduce a third fluid, namely that associated with the kinetic energy of our matter fields, with corresponding energy density $\rho_{\rm KE}$, abundance $\Omega_{\rm KE}$, and equation-of-state parameter $w_{\rm KE}$. It is important to note that we cannot consider $\rho_{M}$ and $\rho_{\rm KE}$ as different contributors to a single common fluid associated with our gas of $\phi$ particles with a unique time-independent equation of state of the form $P=w\rho$ with a constant $w$, since (as we shall demonstrate below) the energy densities associated with rest mass and kinetic energy have different equations of state of this form, with $w_{M}\not=w_{\rm KE}$.
@@ III Dynamical Equations
In general, following Ref. Dienes et al. (2024a), we shall refer to any process that induces a transfer of energy between two energy components as a “pump.” Such pumps are associated with local particle-physics processes which conserve energy density, but redistribute it among our different energy components.
@@ IV Pump Terms for Thermal Annihilation
It is not difficult to obtain explicit expressions for our pump terms. Within the cosmology we are studying, such pumps represent the effects of the two-body annihilations of $\phi$ particles into radiation. Such process are familiar from traditional studies of the thermal freeze-out phenomenon, where one has $dn_{M}/dt\sim\langle\sigma v\rangle n_{M}^{2}+...$ where $n_{M}$ is the matter-particle number density. In this expression, $\langle\sigma v\rangle$ denotes the thermally averaged “swept-volume” rate, i.e., the rate at which volume is swept by the cross-section $\sigma$ moving with transverse velocity $v$. Recognizing that $n_{M}=\rho_{M}/m$, we see that our corresponding pump is nothing but
@@
where $|\vec{p}_{\rm CM}|\equiv|\vec{p}_{a}-\vec{p}_{b}|$ is the magnitude of the three-momentum of either of the two incoming particles $\phi_{a}$ and $\phi_{b}$ in the CM frame, where $q$ is an arbitrary exponent, and where $C$ is a momentum-independent and $q$-independent coefficient. The ansatz in Eq. (16) is not entirely unfamiliar; for example, a similar ansatz is often invoked when discussing thermal freeze-out (see, e.g., Ref. Kolb and Turner (1990); Jungman et al. (1996); Bertone et al. (2005)), with the exponent $q$ selecting between $s$-wave annihilation ($q=0$), $p$-wave annihilation ($q=2$), $d$-wave annihilation ($q=4$), and so forth. However, although the value of $q$ naïvely corresponds to twice the order of the leading term in the partial-wave expansion, $q$ can actually be more negative than stated above if our annihilation proceeds through an $s$-channel process with a propagator that is in resonance at low momenta. Furthermore, Sommerfeld enhancement Sommerfeld (1931) can also give rise to negative values of $q$. For this reason, we will also allow ourselves to consider values of $q$ which are negative. Indeed, we shall give an example of a model that satisfies Eq. (16) with such values of $q$ in Sect. VIII.
@@ V Dynamical Equations Redux and the Role of Coldness
Given these pumps, we can now construct the differential equations that govern the time-evolution of our system within this cosmology. In principle, there are two quantities whose variations are of interest to us: $\Omega_{M}$ and $\Omega_{\rm KE}$. Indeed, as is evident from Eq. (2), the temperature $T$ is simply related to the quotient $\Omega_{\rm KE}/\Omega_{M}$. However, $\rho_{M}$ remains as an additional independent quantity within this system because the expressions in Eq. (25) for our energy-density pumps depend quadratically on $\rho_{M}$. Indeed, although one of these factors of $\rho_{M}$ is converted to $\Omega_{M}$ when constructing our abundance pumps $P_{ij}$ in Eq. (10), the other factor of $\rho_{M}$ remains unconverted. Thus, in this system, our differential equations cannot be reduced to abundances alone, and the absolute matter energy density $\rho_{M}$ — or equivalently the Hubble parameter — remains as an additional degree of freedom. This is a new feature that did not appear in any of the previous stasis analyses in Refs. Dienes et al. (2022a, b, 2024a). However, this feature arises in the present case as a new consequence of our choice of pump.
@@ VI Fixed-Point Solutions and Thermal Stasis
Interestingly, we see that the fixed-point abundance ${\overline{\Omega}}_{M}$ does not depend on the pump prefactor $\widehat{C}$. This is ultimately the case because the dynamical equations in Eq. (28) are invariant under the simultaneous transformations $\Xi\to\alpha\Xi$ and $\widehat{C}\to\widehat{C}/\sqrt{\alpha}$, where $\alpha$ is an arbitrary scaling parameter. This observation is analogous to the observation that the quantity ${\overline{\Omega}}_{M}$ in Ref. Dienes et al. (2022a) is independent of $\Gamma_{0}$, as well as similar observations in Refs. Dienes et al. (2022b, 2024a).
@@
We thus conclude that our system necessarily evolves into an extended stasis epoch, regardless of its initial conditions pertaining to either the initial matter (or radiation) abundance or the initial coldness. This is therefore a direct demonstration that the stasis phenomenon can be extended into the thermal domain, with a prolonged period of stasis experienced not only by the energy abundances but also by thermodynamic quantities — such as coldness — which involve the temperature of the matter prior to annihilation.
@@
Given this state of affairs, it is natural to wonder how this stasis is able to persist across so many $e$-folds of cosmological expansion — a concern heightened by the fact that only a single matter species is involved. Ultimately, this feature is tied to the unique form of our pump. As we have seen, our $\phi$-particle annihilation pumps for both energy density and kinetic-energy density scale as
@@ VII Thermal Stasis as a Global Attractor
It is straightforward to demonstrate that the non-trivial fixed-point solutions for $({\overline{\Omega}}_{M},{\overline{\Xi}})$ in Eq. (32) are also global attractors. We have already seen evidence of this attractor behavior in Fig. 2. As a first step, we note that for a given value of $q$, we may assess whether the corresponding fixed-point solution is a local attractor by evaluating the $2\times 2$ Jacobian matrix for our dynamical equations for $d\Omega_{M}/d{\cal N}$ and $d\Xi/d{\cal N}$ and then examining the signs of its two eigenvalues $\lambda_{\pm}$ at that fixed point. In Fig. 3, we present a parametric plot of these eigenvalues as $q$ is varied within the range $q_{\rm min}<q<q_{\rm max}$ for which our non-trivial fixed-point solution exists. These eigenvalues trace out the red solid curve. Within this $q$-range, we see that both eigenvalues lie within the yellow-shaded region wherein both $\lambda_{+}$ and $\lambda_{-}$ are negative. Thus, we may conclude that across this entire range of $q$, the non-trivial fixed-point solution within this range is indeed a local attractor.